Four cards are drawn from a desk without replacement. Find these probabilities.




a. All cards are jacks.




b. All cards are black cards.




c. All cards are hearts.





1
Expert's answer
2022-06-09T07:23:00-0400

The probability formula states:

"P(A)=\frac{\text{number of favorable events}}{\text{number of total events}}."


There are 52 cards in a deck, so there are

"C_{52}^4=\frac{52!}{4!\cdot(52-4)!}=\frac{52}{4!48!}=\frac{52\cdot51\cdot50\cdot49}{4\cdot3\cdot2\cdot1}=270,725"

ways for getting 4 cards.


a. There are 4 jacks in a deck, so there is

"C_4^4=\frac{4!}{4!\cdot(4-4)!}=\frac{4!}{4!0!}=1"

way to get all of them.

"P(\text{All cards are jacks})=\frac{C_4^4}{C_{52}^4}=\frac{1}{270,725}\approx" 0.0000037.


b. There are 26 black cards in a deck (13 clubs and 13 spades), so there are

"C_{13}^4=\frac{26!}{4!\cdot(26-4)!}=\frac{26}{4!22!}=\frac{26\cdot25\cdot24\cdot23}{4\cdot3\cdot2\cdot1}=14,950"

ways to get 4 of them.

"P(\text{All cards are black cards})=\frac{C_{26}^4}{C_{52}^4}=\frac{14,950}{270,725}\approx" 0.055222.


c. There are 13 hearts in a deck, so there are

"C_{13}^4=\frac{13!}{4!\cdot(13-4)!}=\frac{13!}{4!9!}=\frac{13\cdot12\cdot11\cdot10}{4\cdot3\cdot2\cdot1}=715"

ways to get 4 of them.

"P(\text{All cards are hearts})=\frac{C_{13}^4}{C_{52}^4}=\frac{715}{270,725}\approx" 0.002641.


Answer. a. "\frac{C_4^4}{C_{52}^4}\approx" 0.0000037

b. "\frac{C_{26}^4}{C_{52}^4}\approx" 0.055222

c. "\frac{C_{13}^4}{C_{52}^4}\approx" 0.002641